OpenStudy (anonymous):

How to find the derivative of y=x^(cosx)? Reeeally need help with this.. Have an exam in 2 hours...

OpenStudy (anonymous):

Logarithmic differentiation would help. Write, \[y=x^{\cos x}~~\iff~~\ln y=\ln\left(x^{\cos x}\right)\] Log properties tell you that \[\ln y=\cos x\ln x\] Then differentiate both sides, which will involve chain and product rules: \[\frac{d}{dx}\ln y=\frac{d}{dx}\cos x\ln x\\ \frac{1}{y}\frac{dy}{dx}=-\sin x\ln x+\cos x\frac{1}{x}\] Then solve for \(\dfrac{dy}{dx}\).

OpenStudy (anonymous):

THANK YOU THAT MAKES SO MUCH SENSE.

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):